OpenStudy (mattyboyy):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0?

OpenStudy (anonymous):

rewrite the equation in this form y=mx+b

OpenStudy (mattyboyy):

How? I honestly forgot.

OpenStudy (mattyboyy):

Combine like terms?

OpenStudy (anonymous):

3(y − 2) + 6(x + 1) − 2 = 0? 3y-6+6x+6-2=0 3y+6x=2 3y=-6x+2 y=-2x+3/2

OpenStudy (mattyboyy):

Thanks thats what I got

OpenStudy (anonymous):

that is good

OpenStudy (anonymous):

so do you know the slope and y intercept

OpenStudy (mattyboyy):

Yes the slope is -2 and the y is 3/2 thanks